OpenStudy (anonymous):

What is the equation of the line that has a slope of 4 and passes through the point (3,-10) Please include steps :) Thanks

jimthompson5910 (jim_thompson5910):

y = mx+b -10 = (4)(3)+b -10 = 12+b -10-12 = b -22 = b b = -22 So the equation of the line is y = 4x-22

OpenStudy (anonymous):

Thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome